OpenStudy (anonymous):

27 is what percent of 9?

OpenStudy (anonymous):

ok this is pretty easy i will explain it to you whenever you have a percent and you want to know the percent of a number then you take your percent which in this case is 27% and multiply it my the number which is 9 and to multiply a percent my a number you take the percent (27%) and make it a decimal so to make a percent a decimal you take have to move the decimal but the decimal is invisible here it is right inbetween the 7 and the % so move the invisible decimal over 2 spaces so that gives you .27 and now you take 0.27 and multiply it my 9 so.. 0.27(9)=2.43 and that is how you do it i hope it helped you and please mark this as helpful :)

OpenStudy (anonymous):

Don't forget, OF means * and is means =

OpenStudy (anonymous):

27 should be larger than 9, so 100% of 9=9. 27 is larger than that. in algebraic form, to solve this you'd write 9 * x/100 = 27 9x/100=27 9x= 27*100 9x=2700 x=2700/9 x=300 so, 27 is 300% of 9, and not 2.43, which is 27% of 9.